Loli:

Which expression is equivalent to (5y^2)(2y^-3) A~ 7y^-6 B~ 10y^-6 C~ 7y^-1 D~ 10y^-1

Vocaloid:

start by adding the exponents

Vocaloid:

2 - 3 = ?

Loli:

2 - 3 = -1

Vocaloid:

good so our y term becomes y^(-1)

Vocaloid:

then we just multiply the constants 5*2 = 10

Vocaloid:

giving us 10y^(-1)